DeMarcus wants to retire with​ $1 million in savings by the time he turns 70. He is currently 18 years old. How much will he need to save each​ year, assuming he can get a 10​% annual return on his​ investments?
Answer #1

Using financial calculator input

FV=1000000

N= 70-18= 52

I/Y=10

Solve for PMT as -709

Hence annual saving = $709